AO SDT on 07 November 2011 UT

Observer(s):C. Arcidiacono, F. Manucci, J. Argamedo, E. Pinna, A.Skemer,JC. Guerra (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: T. Sargent (Tucson)

Telescope Support: M. Midkiff (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Dome not open. Lost the night.
Not observing due to bad weather and ice on the top. 
Humidity:98%
Temperature outside:-7
Retro-reflector ON.

Test of close loop with the retro-reflector. 

Details

Test with the observe routines.

Systems warm up (W-unit and secondary), secondary went to power off, the W-unit power off. Telescope chiller power was off. David power on the chillers. The W-unit has stopped a few times due to failure of the ts847. Once the CCD47 restarted the procedures (acquire ref, or close loop ) went forward.. Checking stability of the close loop

-- JuanGuerra - 06 Nov 2011
Topic revision: r6 - 13 Nov 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback